### v2

# Removing singularities for fully nonlinear PDEs

Abstract
We show removability of half-line singularities for viscosity solutions of fully nonlinear elliptic PDEs which have classical density and a Jacobi inequality. An example of such a PDE is the Monge-Ampère equation, and the original proof follows from Caffarelli 1990. Other examples are the minimal surface and special Lagrangian equations. The present paper’s quick doubling proof combines Savin’s small perturbation theorem with the Jacobi inequality. The method more generally removes singularities satisfying the single side condition.
1 Introduction
1.1 Monge-Ampère
We establish a new proof of a known result that the Monge-Ampère equation
$F(D^{2}u)=\det D^{2}u=1,\qquad x\in B_{1}(0)\subset\mathbb{R}^{n}$ (1.1)
does not have viscosity solutions singular inside a half-line. Denote by $\text{sing}(u)\subset B_{1}(0)$ the relative complement of those points on which $u$ is twice differentiable.
Theorem 1.1.
Let $u\in C(\overline{B_{1}(0)})$ be a convex viscosity solution of (1.1). If $\text{sing}(u)$ is contained in a closed line segment with an endpoint in $B_{1}(0)$, then $\text{sing}(u)=\varnothing$.
Remark 1.1 (Singularities of Monge-Ampère).
This result originally follows from the regularity theory of Monge-Ampère. The singular set is well known by Pogorelov’s Hessian estimate [P64, P78] to be the set of points where $u$ fails strict convexity, and is closed in $B_{1}(0)$. It follows from Caffarelli [C90] that $u$ only equals a linear function on sets with no extremal points: subspaces. For example, the Pogorelov [P78] singular solution $u(x)=|(x_{1},x_{2})|^{1+\alpha}f(x_{3})$ is singular along a full line. The dimension of degeneration is less than $n/2$, as follows from Caffarelli [C93], where examples of such dimensions were also produced. It follows from Mooney [M15] that the collection of $k$-dimensional degenerate subspaces has Hausdorff dimension less than $n-k$, and that $\text{sing}(u)$ has sharp upper bound of dimension $n-1$. In dimension two, Alexandrov showed strict convexity [A42]. Singular set propagation to the boundary was obtained for Gaussian curvature and curvature functions by Smith [S12, S24]. Namely, the singular set is the union of the convex hull of closed subsets of the domain boundary.
Remark 1.2 (Improvability).
Our result can remove half-line singularities, which seems optimal in dimension three. However, in higher dimensions, the singularities are codimension two planes. Can one remove half-plane singularities using a similar proof? This does not seem clear using the elementary methods of this paper.
1.2 General PDEs: minimal surface, special Lagrangian equation
Our next result extends Theorem 1.1 to fully nonlinear equations which are classically dense and have a Jacobi inequality, which includes the minimal surface equation
$D\cdot\left(\frac{Du}{\sqrt{1+|Du|^{2}}}\right)=0,$ (1.2)
and the special Lagrangian equation
$\text{tr}\arctan D^{2}u=c\in[(n-2)\pi/2,n\pi/2).$ (1.3)
We encapsulate these equations and Monge-Ampère in the following definition.
Definition 1.1.
Let $F:\text{Sym}(n;\mathbb{R})\times\mathbb{R}^{n}\to\mathbb{R}$ be Lipschitz and elliptic $F(M+P,p)\geq F(M,p)$ for $P$ positive definite. We assume the equation is smooth and strictly elliptic at every point of a smooth solution, or $\partial F(M,p)/\partial M>0$ at all points $(M,p)=(D^{2}u(x),Du(x))$. The fully nonlinear equation
$F(D^{2}u,Du)=0,\qquad x\in B_{1}(0)\subset\mathbb{R}^{n}$ (1.4)
has classical density if for every viscosity solution $u\in C(\overline{B_{1}(0)})$, there exists a sequence $u_{k}\in C^{\infty}(\overline{B_{1}(0)})$ of smooth solutions to (1.4) converging uniformly to $u$.
It is said to have a Jacobi inequality if there exists a Lipschitz positive function $b(M,p)$ satisfying the Jacobi inequality on all such smooth solutions $u$ in the viscosity sense,
$F^{ij}b_{ij}$ $\geq 2F^{ij}b_{i}b_{j}/b,$ (1.5)
$F^{ij}$ $:=\partial F/\partial M_{ij}(D^{2}u,Du),$
$b_{i}$ $:=\partial b(D^{2}u,Du)/\partial x^{i},$
and the property that $b(D^{2}u,Du)\leq B$ on a domain $\Omega\subset\subset B_{1}(0)$ implies higher bounds
$|D^{k}u|\leq C(k,B,\Omega^{\prime},\Omega,\|u\|_{L^{\infty}(B_{1}(0))}),\qquad k=0,1,2,3\dots{}$ UNKNOWN (1.6)
on any subdomain $\Omega^{\prime}\subset\subset\Omega$ and smooth solution $u$.
Theorem 1.2.
Let $u\in C(\overline{B_{1}(0)})$ be a viscosity solution of equation (1.4) having classical density and a Jacobi inequality. If $\text{sing}(u)$ is contained in a line segment with an endpoint in $B_{1}(0)$, then $\text{sing}(u)=\varnothing$.
Remark 1.3 (New proof for minimal surface).
The Jacobi inequality for $b=\sqrt{1+|Du|^{2}}$ is true for minimal surface (1.2), see e.g. [GT77, Eq. (16.49)]. The higher bounds (1.6) follow from the $C^{0,1}\to C^{1,\alpha}$ local estimate of De-Giorgi Nash. See e.g. [GT77, Theorem 16.8] for classical solvability of the Dirichlet problem for smooth solutions, then classical density by the comparison principle for viscosity solutions by smoothly approximating the boundary data. This gives a new proof of half-line singularity removal; the original proof is by interior regularity, which follows from the gradient estimate of Bombieri-De Giorgi-Miranda [BDM69], see also Trudinger [T72].
Remark 1.4 (New proof for special Lagrangian equation).
The Jacobi inequality for $b=(1+\lambda_{max}(D^{2}u)^{2})^{\varepsilon(n)/2}$ and $n\geq 3$ is true for special Lagrangian equation by the proof of [WY14, Proposition 2.1]. The higher bounds (1.6) follow follow from the local estimates of Evans-Krylov $C^{1,1}\to C^{2,\alpha}$, using the level set convexity in [Y06]. The classical density follows from the classical solvability in Caffarelli-Nirenberg-Spruck [CNS85]; see also the lecture notes of Yuan [Y10] and the ones [Y04, B20] for the comparison principle for viscosity solutions. This new proof of half-line singularity removal comes after the interior regularity; see [WY09, WY14] for Hessian estimates and [WY10] for gradient estimates.
Remark 1.5 (New proof for Monge-Ampère equation).
We show that Theorem 1.1 follows from Theorem 1.2. Let us first note the standard facts that operator $F(D^{2}u)$ can be defined elliptic everywhere by extending $F=\det D^{2}u-1$ to $-1$ outside the positive definite matrices, and viscosity solutions are defined by touching $u$ from above and below by convex polynomials. The comparison principle for viscosity solutions, using again the general proof written down in e.g. [Y04, B20], works verbatim for the Monge-Ampère equation.
Next, the Monge-Ampère equation has a Jacobi inequality [Y23] for $b=\det(I+D^{2}u)^{1/2n}$. By Evans-Krylov, the higher estimates (1.6) work. Finally, the classical solvability [CNS84, K84] and comparison principle for Monge-Ampère implies the classical density. Thus singularity removal Theorem 1.1 works.
Remark 1.6 (Other PDEs with Jacobi inequalities).
There are many other nonlinear PDEs which have Jacobi or Jacobi type inequalities, at least under some a priori conditions on the solutions. To list a few recent ones, there are the Lagrangian mean curvature equations [B21, L23a, Z23, BW24a, BW24b], generalized special Lagrangian equation [Z24], the Hessian quotient equation [L23b, L24], scalar curvature equations [CJZ24, QZ24, W24], and 2-Hessian equation with dynamic semi-convexity [ShY23].
1.3 Extensions
Our last result extends the removable singularities from half-lines to sets satisfying the single side condition, which roughly states that the singularity only intersects a single side of some box. It excludes singularities which intersect two sides of any closed box. Half-lines and point singularities are basic examples of single side singularities, while full lines and half-planes fail this condition.
We recall that a closed box in $B_{1}(0)$ is a rotation about the origin of a set of the form
$\{x\in\mathbb{R}^{n}:a_{i}\leq x_{i}\leq b_{i},1\leq i\leq n\},\qquad a_{i}<b_{i}\quad\text{for all i},$
which is contained in $B_{1}(0)$. The sides of the box are closed pieces of hyperplanes which comprise its boundary. We also denote $B^{0}$ by the interior of a box $B$.
The theorem states that the restriction of a singularity to a box is removable if it only intersects a single side of the box.
Theorem 1.3 (Monge-Ampère).
Let $u\in C(\overline{B_{1}(0)})$ be a convex viscosity solution of (1.1) on domain $B_{1}(0)\subset\mathbb{R}^{n}$. If there exists a closed box $B$ in $B_{1}(0)$ such that $\text{sing}(u)$ intersects only one side of $B$, then $\text{sing}(u)\cap B^{0}=\varnothing$.
Theorem 1.4 (General PDEs).
Let $u\in C(\overline{B_{1}(0)})$ be a viscosity solution of (1.4) having classical density and Jacobi inequality (1.5). If there exists a closed box $B$ in $B_{1}(0)$ such that $\text{sing}(u)$ intersects only one side of $B$, then $\text{sing}(u)\cap B^{0}=\varnothing$.
Example 1.1 (Keyhole singularities).
By iterating this result, we can remove singularities contained in a keyhole, i.e. a pyramid inside $\Omega$ for which a single vertex intersects the boundary. Moreover, there are removable singularities not contained in a keyhole. One can remove any singularity whose closure intersects the boundary only once.
1.4 Discussion
Our results show that the singularities of solutions of a PDE must obey geometric constraints if the PDE has a Jacobi inequality. Moreover, the Pogorelov singular solution of the Monge-Ampère equation (1.1) shows that the geometric condition in Theorem 1.3 is rather sharp. Whether it can be improved is unclear to us; see Remark 1.2.
An earlier, standard result is the removability of point singularities for $C^{\infty}$-solvable equations, i.e. those PDEs which have classical solutions to Dirichlet problems with smooth boundary data, such as concave PDEs as in [CNS85]. A viscosity solution has a point singularity at $p\in\Omega$ if $u\in C^{\infty}(\overline{\Omega}/\setminus\{p\})\cap C^{0}(\overline{\Omega})$. The result states that if a $C^{\infty}$-solvable equation has a solution with a point singularity, then it is smooth. The proof is by the comparison principle.
In contrast, removing singularities which intersect the boundary is not well understood. The global estimates break if the boundary values can be singular. Our results show that if the PDE has a Jacobi inequality, the singularities can be removed even if the boundary values are singular. The only requirement is that the intersection of the singularity with the boundary be sufficiently small. For example, a half-line intersects the boundary only at a point. More general removable singularities satisfy the single side condition, which enforces a weak transversality of the singularity to the boundary.
Other results on the partial regularity for fully nonlinear PDEs include sharp Hausdorff dimension upper bounds of $n-\varepsilon$ in [ASS12] for uniformly elliptic PDEs, and $n-1$ by Mooney [M15] for the Monge-Ampère equation. Applying Savin [S07] to PDEs with an Alexandrov theorem, such as for convex functions or k-convex functions and solutions, see [CT05] and [ShY23, Remark 5.2], shows that the singular set is closed and Lebesgue measure zero.
1.5 Outline of the doubling proof
Our main observation is a doubling inequality (2.4) which controls $b(D^{2}u,Du)$ inside $[-1,0]\times[-1,1]$, for example, by its values on three sides of this rectangle.
A doubling inequality can be thought of as a maximum principle that fails on a small set. It controls the quantity of interest by its values on that set. Given a Jacobi inequality, or equivalently a positive superharmonic function, a doubling inequality can often be proved. One proof is by a comparison argument with a Green’s function, as in [SY23, Lemma 2.6, pg 4056]. Another proof is by Korevaar type arguments [S24, SY24]. An integral doubling inequality was shown early by Trudinger [T80, pg 70] using the ABP estimate. Using an almost-Jacobi inequality, for which a superharmonic function is not available, the doubling inequality of Qiu [Q24] was extended in [ShY23] using the comparison function of Guan-Qiu [GQ19].
Our approach follows [SY23] and compares with a “1D Green’s function” $(1/|x_{1}|-1)_{+}$. In [SY23], the Green’s function vanishes on the boundary, so the defective set is near the pole of the Green’s function. But in our situation, the Green’s function is nonvanishing near $x_{2}=\pm 1$ as well as the pole $x_{1}=0$. The doubling inequality only controls $b(D^{2}u,Du)$ in terms of its values on these sides of a box. But provided the singular set, such as a half-line, satisfies the single side condition, it follows that $b(D^{2}u,Du)$ is bounded on the three bad sides, and we obtain a global estimate.
The final ingredient in the proof is Savin’s small perturbation theorem [S07, Theorem 1.3], which lets us link the a priori estimate to the partial regularity assumption away from the singular set. If we use the smooth approximations from classical density, the smooth solutions will be uniformly bounded on the smooth set. Applying this boundedness to the a priori doubling inequality gives global higher derivative bounds on the smooth approximations. The limit will therefore be interior smooth.
2 Proof of Theorems
We prove Theorem 1.4, since the others follow from this one.
Step 1. Single side singularity. Let $u$ be a viscosity solution of (1.4), and suppose there exists a closed box $B$ in $B_{1}(0)$ such that $\text{sing}(u)\cap\partial B$ is contained in the interior of one of the sides of $\partial B$. After a rotation, we write this box as as a set of the form $\{a_{i}\leq x_{i}\leq b_{i}\}$, and we suppose $\text{sing}(u)\cap\partial B\subset L^{0}$ for left side $L:=\{x_{1}=a_{1}\}\cap B$. Since $\text{sing}(u)$ is closed, we can find a slightly smaller box $B^{\prime}$ satisfying the same hypotheses by replacing $b_{1}$ with $b_{1}^{\prime}=b_{1}-r$ for some $r>0$.
Now let us denote $R:=\partial B^{\prime}\setminus L^{0}$ as the “right” portion of $\partial B^{\prime}$ which excludes the left side $L$. The importance of the single side condition is that closed set $\text{sing}(u)$ has a positive distance from $R$. Indeed, this follows from closed set $\text{sing}(u)\cap\partial B$ being contained in open set $L^{0}$. Thus, $R$ is compactly contained in the smooth, open set $\text{sing}(u)^{c}=B_{1}(0)\setminus\text{sing}(u)$.
Step 2. Doubling inequality on smooth approximations. By classical density, we find smooth solutions $u_{k}\in C^{\infty}(\overline{B_{1}(0)})$ solving (1.4) which converge to $u$ uniformly as $k\to\infty$. Letting $b_{k}=b(D^{2}u_{k},Du_{k})$, we compare with a “1D Green’s function” and form the test function on box $B^{\prime}$
$w(x)=\eta(x_{1})b_{k}(x),\qquad\eta(x_{1})=\frac{1}{b_{1}-x_{1}}-\frac{1}{b_{1}-a_{1}}.$ (2.1)
Suppose $w$ achieves its maximum value $W>0$ inside $B^{\prime}$ at $x=p$, or $w(x)\leq W$ with equality at $x=p$. Then $b_{k}(x)\leq W/\eta(x_{1})$ with equality at $x=p$, i.e. $W/\eta$ touches $b_{k}$ from above at $x=p$. Since the Jacobi inequality (1.5) is true in the viscosity sense, it follows that, at $x=p$,
$F^{11}\left(\frac{1}{\eta}\right)^{\prime\prime}\geq 2F^{11}\eta^{\prime}(x_{1})^{2}/\eta^{3}.$ (2.2)
Since $F^{11}>0$ by ellipticity on smooth solutions, this simplifies to
$0>-\eta^{\prime\prime}/\eta^{2}\geq 0$ (2.3)
This contradiction implies the maximum occurs on the boundary. Recalling right portion $R$ in Step 1, we obtain
$\sup_{B^{\prime}}\eta(x_{1})b_{k}(x)$ $=\sup_{\partial B^{\prime}}\eta(x_{1})b_{k}(x)$ (2.4)
$\leq r^{-1}\sup_{R}b_{k}(x).$
Thus we obtain control of $b_{k}$ by its values on a piece of the boundary.
Step 3. Savin stability of partial regularity on $R$. By uniform convergence $u_{k}\to u$ in $B_{1}(0)$ and Savin [S07, Theorem 1.3], it follows that $u_{k}\to u$ locally smoothly in the smooth set, or in $C^{2,\alpha}_{loc}(\text{sing}(u)^{c})$. On the other hand, by Step 1, $R$ is compactly contained in $\text{sing}(u)^{c}$. It follows that $Du_{k}(x)$ and $D^{2}u_{k}(x)$ are uniformly bounded on $R$ for all $x$ and $k$. Therefore, there exists $A>0$ such that, for all $k$,
$\sup_{B^{\prime}}\eta(x_{1})b_{k}(x)\leq A.$ (2.5)
Step 4. Interior regularity on $B^{0}$. Fix $x_{0}\in(B^{\prime})^{0}$. For any $x\in(B^{\prime})^{0}$, we find that $b(D^{2}u_{k}(x),Du_{k}(x))=b_{k}(x)\leq C(x_{1},A)$ is locally bounded uniformly in $k$, with blowup only as $x_{1}\to a_{1}$. By the higher local bounds (1.6) and eventual boundedness of $\|u_{k}\|_{L^{\infty}(B_{1})}$, it follows that $|D^{m}u_{k}(x)|\leq C(m,A,x_{1},\|u\|_{\infty})$ is bounded nearby $x_{0}$. After a subsequence, we assume $u_{k}\to u$ in $C^{2,\alpha}$ nearby $x_{0}$. We conclude $u$ is $C^{2,\alpha}$ nearby $x_{0}$. Varying $x_{0}$, this shows $\text{sing}(u)\cap(B^{\prime})^{0}=\varnothing$, hence $\text{sing}(u)\cap B^{0}=\varnothing$. This completes the proof.
3 Appendix
Savin small perturbation theorem. We specialize Savin [S07, Theorem 1.3] to equations $F(M,p,x)=0$ defined on $\text{Sym}(n;\mathbb{R})\times\mathbb{R}^{n}\times B_{1}(0)$, for Definition 1.1.
Proposition 3.1 ([S07], Theorem 1.3).
Let $F(D^{2}u,Du,x)=0$ satisfy the following hypotheses:
(i) $F$ is locally Lipschitz and elliptic, or $F(M+N,p,x)\geq F(M,p,x)$ if $N>0$, with strict ellipticity in a neighborhood of $M=p=0$, $\partial F/\partial M>0$,
(ii) $F(0,0,x)=0$ for all $x$,
(iii) $F\in C^{2}$ in a neighborhood of $M=p=0$.
Then there exists $c_{1}$ small enough depending on $n,F$ such that if viscosity solution $u$ of $F(D^{2}u,Du,x)=0$ satisfies flatness $\|u\|_{L^{\infty}(B_{1}(0))}\leq c_{1}$, then $u\in C^{2,\alpha}(B_{1/2})$ with $\|u\|_{C^{2,\alpha}(B_{1/2})}\leq 1$.
Savin stability of partial regularity. Suppose $u$ is a viscosity solution of elliptic PDE $G(D^{2}u,Du)=0$ satisfying the assumptions of Definition 1.1, which is smooth on a subdomain $\Omega\subset B_{1}(0)$, with smooth solutions $u_{k}\to u$ uniformly on $B_{1}(0)$. Then by Savin, $u_{k}\to u$ in $C^{k}_{loc}(\Omega)$ for any $k$, i.e. $C^{k}$ convergence in any compact subset of $\Omega$. Indeed, let $v_{k}=u_{k}-u$ solve the elliptic equation on $\Omega$
$F(M,p,x)=G(D^{2}u(x)+M,Du(x)+p)-G(D^{2}u(x),Du(x))$
which is $C^{\infty}$ near points $M,p=0$ and satisfies Savin’s assumptions. Since $\|v_{k}\|_{\infty}\to 0$ as $k\to\infty$, it follows from Savin and a covering argument on small enough balls that $v_{k}$ has locally uniform-in-$k$ $C^{2,\alpha}$ and higher estimates as $k\to\infty$.